# Quantum fluctuation energies over a spatially inhomogeneous field background in a chiral soliton model

### v3

@@ I Introduction
In recent years there are lots of investigations on the possible inhomogeneous phases in quantum chromodynamics (QCD) systems like the chiral density wave(CDW) phase, the chiral spiral phase, the chiral soliton lattice(CSL) phase and etc.Chernodub (2021); Chen et al. (2022); Fukushima and Sasaki (2013); Buballa and Carignano (2015); Brauner and Yamamoto (2017); Nickel (2009). These studies have shown extensive interests in the theoretical exploration about the novel ground states or new phase boundaries in quark/nuclear matterIshioka et al. (2010); Evans et al. (2012); Voskresensky (2023); Lakaschus et al. (2021); Endrődi et al. (2021). Due to recent advancements in studies of the field of rotating neutron stars and in-depth experimental and theoretical exploration of quark matter at finite densities, the possibility of inhomogeneous phases in quark matter has been considered more seriouslyCardall et al. (2001); Tsokaros et al. (2022); Bzdak et al. (2020); Ferrer et al. (2024). The existence of such inhomogeneous phases is theoretically a semi-classical field configuration formed due to strong coupling effects, resulting in a bound state that creates a spatially inhomogeneous field background, thereby leading to new phases of quark matter on this background. Typically, these semi-classical backgrounds are calculated based on the mean field approximation. However, in the usual discussions of inhomogeneous phases the quantum fluctuation effects have not been systematically considered or rigorously calculated. The calculation of quantum fluctuations on inhomogeneous field backgrounds is very complicated, and even defining a consistent calculation framework within field theory is challenging. Previous studies have used derivative expansion methods, but these studies are given under certain approximations, namely when the spatial variation of the field is relatively slowLi and Perry (1988); Stewart and Blunden (1997). The use of mode truncation methods in finite volumes and ultraviolet truncation methods brings corresponding uncertaintiesKahana and Ripka (1984). In supersymmetric models, the study of quantum corrections in topological soliton background is of particular interest. Due to the supersymmetry, the energy of the soliton will satisfy the Bogomol’nyi-Prasad-Sommerfield (BPS) relation which indicates there is a lower bound of the energy. At the classical level this bound is saturated. At quantum level, since the quantum corrections to the energy and the central charge precisely counterbalance each other, the BPS relations are still saturated to equalities. However, the special results rely on topological soliton background and supersymmetryShifman et al. (1999); Dunne (1999); Graham and Jaffe (1999); Rebhan et al. (2004, 2005). Farhi, Graham and Jaffe, among others, have developed a spectral method that can consistently solve for quantum corrections on general soliton backgroundsFarhi et al. (1998, 2002); Graham et al. (2009). This method is self-consistent and does not introduce additional uncertainties, allowing for the study of quantum fluctuations of topological and non-topological solitons. In this method due to the inhomogeneous background the usual density of states of the momentum in momentum integration of the standard one loop quantum fluctuation energy will receive a correction which can be derived by solving the field equations and evaluating the scattering phase shift in the inhomogeneous background. This method was first proposed by Schwinger in calculations of the electron’s quantum fluctuations in certain nontrivial configurations of electromagnetic fields as the backgroundSchwinger (1954), and later used by ‘t Hooft to calculate the quantum corrections of instantonst Hooft (1976), and later developed by Farhi, Graham and Jaffe into an effective method for calculating soliton quantum fluctuations.
@@ III.1 One loop quantum fluctuation energy and renormalization
where $d=4-2\epsilon$, $m_{s}=\sqrt{\lambda}f_{\pi}$, and $\mu$ is the scale parameter introduced to make the coupling constant $g$ dimensionless. The effective action can be defined as
@@
Since the asymptotic behavior of $\delta_{G,\omega,\Pi}(k)$ as $k\rightarrow\infty$ is $\frac{1}{k}$, it is the logarithmic divergence. Additionally there are the sums over $G$, $\omega$ and $\Pi$, thus $E_{\rm vac}^{\psi}$ is highly divergent. In order to make it finite, one should adopt a systematic renormalization scheme. Here we have been following the way used in the literatures Graham et al. (2009); Farhi et al. (1998, 2002). As for high momentum, the Born expansion becomes a good approximation to the phase shift. So by subtracting the successive terms in the Born series to a certain order from the phase shift, the divergences could be removed. The regularized scattering phase shift has the following form
@@
where $\delta_{G,\omega,\Pi}^{(\ell)}(k)$ is the $\ell th$ Born approximation to $\delta_{G,\omega,\Pi}(k)$. $\bar{\delta}_{G,\omega,\Pi}(k)$ is also called the subtracted phase shift. However if one subtracts a term in the Born expansion, one needs to add back the equivalent Feynman diagram. The corresponding divergent Feynman diagrams should be renormalized by the counterterms in the usual way Farhi et al. (1998, 2002). Finally one has the finite one loop result of the vacuum fluctuation energy as
@@
where for the subtracted phase shift when $G=0$, the Born subtraction needs to be evaluated to the second order; while for $G\neq 0$ the Born subtraction needs to be evaluated to the fourth order. The phase shifts and Born subtraction of the phase shifts can be calculated by the radial equations, which will be discussed in the next section. The specific calculation of the Born subtraction of the phase shift is presented in Appendix A. According to Ref.Farhi et al. (2002), the adding back divergent Feynman diagrams together with counterterms finally yield the finite pieces $\Gamma_{2}$ and $\Gamma_{4}$ in (40), which forms are
$\Gamma_{2}=$ $\frac{g^{2}}{\pi^{2}}\int_{0}^{\infty}dqq^{2}\left[s^{2}(q)+p^{2}(q)\right]$ (41)
$\times\left\{q^{2}+m_{s}^{2}-6\int_{0}^{1}dx\left[m_{q}^{2}+x(1-x)q^{2}\right]\ln\frac{m_{q}^{2}+x(1-x)q^{2}}{m_{q}^{2}-x(1-x)m_{s}^{2}}\right\}$
[…]
$\left.\left.-2\ln\left(1+x(1-x)\frac{q^{2}}{m_{q}^{2}}\right)\right]\right\},$
$\Gamma_{4}=$ $\frac{g^{4}}{8\pi}\left(\frac{m_{s}^{2}}{m_{q}^{2}}+6\int_{0}^{1}dx\ln\left[1-x(1-x)\frac{m_{s}^{2}}{m_{q}^{2}}\right]\right)$ (42)
$\times\int_{0}^{\infty}drr^{2}\left[\left(\sigma(r)^{2}+\pi(r)^{2}-\sigma_{v}^{2}\right)^{2}-4\sigma_{v}^{2}(\sigma(r)-\sigma_{v})^{2}\right],$
@@ III.2.2 $G\neq 0$ case
where $h_{n}$ is the spherical Hankel function with n=G1,G,G+1n=G-1,G,G+1. In the following we have adopted a strategy which is a little different from that in Farhi’s work Farhi et al. (2002). We do not decouple the first order equations of the upper part and lower part of the Dirac spinor to the second order differential equations. Instead we solve the first order coupled differential equations as a whole. The reason for choosing this strategy is that it is more convenient and efficient in the later calculations of Born subtractions of the phase shifts to higher orders. However, in order to check the results calculated by the first order differential equations, we also carry out the calculations by the second order differential equations which are presented in Appendix B, where the phase shifts and corresponding Born subtraction are also evaluated by the second order differential equations. In Farhi’s work Farhi et al. (2002), the Born subtraction of the phase shift has been only calculated to the second order, while the higher order Born subtraction has been substituted by the fake boson subtraction. In Appendix C, we have made a comparison between the results of the total energies evaluated by the Born subtraction and the fake boson subtraction schemes.
@@
After expanding the first order radial equations to different orders, one can set $g_{s}=g_{p}=1$ in the end. In the expansion we introduce $V^{(n_{s},n_{p})}$, where $n_{s}$ and $n_{p}$ label the order in the expansion around $\sigma_{v}$ and $\pi_{v}$, respectively. The specific calculation process for the Born subtraction of the phase shift by the first order differential equations is detailed in Appendix A, and those calculations for the phase shifts and the corresponding Born subtractions by the second order differential equations are presented in Appendix B.
@@ IV.1 Phase shifts
where $n$ is the number of bound states. Therefore $n=1$ in $G=0$ case means there is one bound state in this case, which is exactly the chiral soliton solution in the ground state. For the cases of $G=0$ and $G=1$ at $\omega=$-, the values of phase shifts begin from $\pi$ at $k=0$, which indicate there are two bound states at negative energy levels for the cases of $G=0$ and $G=1$.
@@
Finally we should mention that the above phase shifts and subtracted phase shifts are evaluated by the first order differential equations. However, we have also checked that all the phase shifts could be evaluated by the second order differential equations as shown in Appendix B, and it gives exactly the same results as those evaluated by the first order differential equations.
@@ IV.2 Quantum fluctuation energies
In table 3 we present the different parts of the energy related to the chiral soliton. One can see that after the summation of $G$, $\omega$ and $\Pi$, the quantum fluctuation energy from the continuum energy spectrum is positive and constitutes the largest part in the vacuum energy, while the quantum fluctuation energy from discrete levels is negative and its magnitude is comparable to that from the continuum spectrum. The renormalization energies from the Feynman diagrams are dominated by $\Gamma_{2}$ which value is negative, while $\Gamma_{4}$ is positive and relatively small. By adding them all one obtains the total renormalized vacuum fluctuation energy over the chiral soliton background, which is negative. If the color degrees of freedom of quarks are taken into account, the vacuum energy should be multiplied by a factor of 3. One could see that the quantum fluctuation energy in magnitude is not small compared to the classical energy of the chiral soliton.
@@ V SUMMARY AND OUTLOOK
Appendix A Calculations of Born subtractions of phase shifts by the first order differential equations

### v4

@@ I Introduction
In recent years there are lots of investigations on the possible inhomogeneous phases in quantum chromodynamics (QCD) systems like the chiral density wave(CDW) phase, the chiral spiral phase, the chiral soliton lattice(CSL) phase and etc.Chernodub (2021); Chen et al. (2022); Fukushima and Sasaki (2013); Buballa and Carignano (2015); Brauner and Yamamoto (2017); Nickel (2009). These studies have shown extensive interests in the theoretical exploration about the novel ground states or new phase boundaries in quark/nuclear matterIshioka et al. (2010); Evans et al. (2012); Voskresensky (2023); Lakaschus et al. (2021); Endrődi et al. (2021). Due to recent advancements in studies of the field of rotating neutron stars and in-depth experimental and theoretical exploration of quark matter at finite densities, the possibility of inhomogeneous phases in quark matter has been considered more seriouslyCardall et al. (2001); Tsokaros et al. (2022); Bzdak et al. (2020); Ferrer et al. (2024). The existence of such inhomogeneous phases is theoretically a semi-classical field configuration formed due to strong coupling effects, resulting in a bound state that creates a spatially inhomogeneous field background, thereby leading to new phases of quark matter on this background. Typically, these semi-classical backgrounds are calculated based on the mean field approximation. However, in the usual discussions of inhomogeneous phases the quantum fluctuation effects have not been systematically considered or rigorously calculated. The calculation of quantum fluctuations on inhomogeneous field backgrounds is very complicated, and even defining a consistent calculation framework within field theory is challenging. Previous studies have used derivative expansion methods, but these studies are given under certain approximations, namely when the spatial variation of the field is relatively slowLi and Perry (1988); Stewart and Blunden (1997). The use of mode truncation methods in finite volumes and ultraviolet truncation methods brings corresponding uncertaintiesKahana and Ripka (1984). In supersymmetric models, the study of quantum corrections in topological soliton background is of particular interest. Due to the supersymmetry, the energy of the soliton will satisfy the Bogomol’nyi-Prasad-Sommerfield (BPS) relation which indicates there is a lower bound of the energy. At the classical level this bound is saturated. At quantum level, since the quantum corrections to the energy and the central charge precisely counterbalance each other, the BPS relations are still saturated to equalities. However, the special results rely on topological soliton background and supersymmetryShifman et al. (1999); Dunne (1999); Graham and Jaffe (1999); Rebhan et al. (2004, 2005). Farhi, Graham and Jaffe, among others, have developed a spectral method that can consistently solve for quantum corrections on general soliton backgrounds Farhi et al. (1998, 2002); Quandt et al. (2017); Graham et al. (2009) . This method is self-consistent and does not introduce additional uncertainties, allowing for the study of quantum fluctuations of topological and non-topological solitons. In this method due to the inhomogeneous background the usual density of states of the momentum in momentum integration of the standard one loop quantum fluctuation energy will receive a correction which can be derived by solving the field equations and evaluating the scattering phase shift in the inhomogeneous background. This method was first proposed by Schwinger in calculations of the electron’s quantum fluctuations in certain nontrivial configurations of electromagnetic fields as the backgroundSchwinger (1954), and later used by ‘t Hooft to calculate the quantum corrections of instantonst Hooft (1976), and later developed by Farhi, Graham and Jaffe into an effective method for calculating soliton quantum fluctuations.
@@ III.1 One loop quantum fluctuation energy and renormalization
where $d=4-2\epsilon$, $m_{s}=\sqrt{\lambda}f_{\pi}$, and $\mu$ is the scale parameter introduced to make the coupling constant $g$ dimensionless Farhi et al. (2002). The effective action can be defined as
@@
Since the asymptotic behavior of $\delta_{G,\omega,\Pi}(k)$ as $k\rightarrow\infty$ is $\frac{1}{k}$, it is the logarithmic divergence. Additionally there are the sums over $G$, $\omega$ and $\Pi$, thus $E_{\rm vac}^{\psi}$ is highly divergent. In order to make it finite, one should adopt a systematic renormalization scheme. Here we have been following the way used in the literatures Graham et al. (2009); Farhi et al. (1998); Quandt et al. (2017); Farhi et al. (2002). As for high momentum, the Born expansion becomes a good approximation to the phase shift. So by subtracting the successive terms in the Born series to a certain order from the phase shift, the divergences could be removed. The regularized scattering phase shift has the following form
@@
where $\delta_{G,\omega,\Pi}^{(\ell)}(k)$ is the $\ell th$ Born approximation to $\delta_{G,\omega,\Pi}(k)$. $\bar{\delta}_{G,\omega,\Pi}(k)$ is also called the subtracted phase shift. However if one subtracts a term in the Born expansion, one needs to add back the equivalent Feynman diagram. The corresponding divergent Feynman diagrams should be properly renormalized Farhi et al. (1998, 2002); Quandt et al. (2017). Finally one has the finite one loop result of the vacuum fluctuation energy as
@@
where for the subtracted phase shift when $G=0$, the Born subtraction needs to be evaluated to the second order; while for $G\neq 0$ case the Born subtraction needs to be evaluated to the fourth order. The phase shifts and Born subtraction of the phase shifts can be calculated by the radial equations, which will be discussed in the next section. $\Gamma_{2}$ and $\Gamma_{4}$ in (40) are finite terms evaluated from the Feynman diagrams. The quadratic divergent fermion diagrams have been renormalized by the corresponding counter terms in the usual way and the result is
$\Gamma_{2}=$ $\frac{g^{2}}{\pi^{2}}\int_{0}^{\infty}dqq^{2}\left[s^{2}(q)+p^{2}(q)\right]\left\{q^{2}+m_{s}^{2}-6\int_{0}^{1}dx\left[m_{q}^{2}+x(1-x)q^{2}\right]\ln\frac{m_{q}^{2}+x(1-x)q^{2}}{m_{q}^{2}-x(1-x)m_{s}^{2}}\right\}$
- $\frac{g^{2}}{\pi^{2}}\int_{0}^{\infty}dqq^{2}p^{2}(q)\left\{m_{s}^{2}+2m_{q}^{2}\int_{0}^{1}dx\left[3\ln\left(1-x(1-x)\frac{m_{s}^{2}}{m_{q}^{2}}\right)\right.\right.\left.\left.-2\ln\left(1+x(1-x)\frac{q^{2}}{m_{q}^{2}}\right)\right]\right\},$ (41)
@@
However, the renormalization of the logarithmic divergence involves a nontrivial calculation. To maintain consistency between the renormalization scheme and the current Born subtraction scheme, we follow the fake boson approach as described in Refs. Farhi et al. (2002); Quandt et al. (2017). The key step is to add back the corresponding logarithmically divergent Feynman diagram and renormalize it. Details of the renormalization procedure using the fake boson method are given in Appendix A. After renormalization, the resulting fake boson Feynman diagram contributes the finite term
$\Gamma_{4}=-c_{B}\int_{0}^{\infty}\frac{dqq^{2}}{2\pi^{2}}\tilde{V}_{B}(q)\tilde{V}_{B}(-q)\left[\int_{0}^{1}dx\frac{x(1-x)}{m_{B}^{2}-x(1-x)m_{f}^{2}}(q^{2}+m_{f}^{2})+\int_{0}^{1}dx\ln\frac{m_{B}^{2}+x(1-x)q^{2}}{m_{B}^{2}-x(1-x)m_{f}^{2}}\right],$ (43)
where
$\tilde{V}_{B}(q)=\int_{0}^{\infty}drr^{2}\frac{\sin qr}{qr}V_{B}(r).$ (44)
$m_{B}$ and $V_{B}(r)$ denote the mass and background potential of the fake boson, respectively. The scale factor $c_{B}$ is determined by matching the logarithmically divergent energy from the third- and fourth-order fermion phase shifts with the corresponding energy from the second-order fake boson phase shift. Meanwhile, $m_{f}$ represents the assigned mass of the external line in the boson loop, which is fixed by matching the renormalization energies in the two fake boson schemes. Further discussions and detailed calculations are provided in Appendix A.
@@ III.2.2 $G\neq 0$ case
where $h_{n}$ is the spherical Hankel function with n=G1,G,G+1n=G-1,G,G+1. In the following we have adopted a strategy which is a little different from that in Farhi’s work Farhi et al. (2002). We do not decouple the first order equations of the upper part and lower part of the Dirac spinor to the second order differential equations. Instead we solve the first order coupled differential equations as a whole. The specific calculation of the Born subtraction of the phase shift by the first order differential equations is presented in Appendix B. The reason for choosing this strategy is that it is more convenient and efficient in the later calculations of Born subtractions of the phase shifts to higher orders. However, in order to check the results calculated by the first order differential equations, we also carry out the calculations by the second order differential equations which are presented in Appendix C, where the phase shifts and corresponding Born subtraction are also evaluated by the second order differential equations. In Farhi’s work Farhi et al. (2002), the Born subtraction of the phase shift has been only calculated to the second order, while the higher order Born subtraction has been substituted by the fake boson subtraction. In Appendix D, we have made a comparison between the results of the total energies evaluated by the Born subtraction and the fake boson subtraction schemes.
@@
After expanding the first order radial equations to different orders, one can set $g_{s}=g_{p}=1$ in the end. In the expansion we introduce $V^{(n_{s},n_{p})}$, where $n_{s}$ and $n_{p}$ label the order in the expansion around $\sigma_{v}$ and $\pi_{v}$, respectively. The specific calculation process for the Born subtraction of the phase shift by the first order differential equations is detailed in Appendix B, and those calculations for the phase shifts and the corresponding Born subtractions by the second order differential equations are presented in Appendix C.
@@ IV.1 Phase shifts
where $n$ is the number of bound states. Therefore $n=1$ in $G=0$ case means there is one bound state in this channel, which is exactly the chiral soliton solution in the ground state. For $G=0$ and $G=1$ at $\omega=$-, the phase shifts start from $\pi$ at the threshould ($k=0$), which indicates the existence of one bound state in each of these channels at negative energies.
@@
Finally we should mention that the above phase shifts and subtracted phase shifts are evaluated by the first order differential equations. However, we have also checked that all the phase shifts could be evaluated by the second order differential equations as shown in Appendix C, and it gives exactly the same results as those evaluated by the first order differential equations.
@@ IV.2 Quantum fluctuation energies
In table 3 we present the different parts of the energy related to the chiral soliton. One can see that after the summation of $G$, $\omega$ and $\Pi$, the quantum fluctuation energy from the continuum energy spectrum is positive and constitutes the largest part in the vacuum energy, while the quantum fluctuation energy from discrete levels is negative and its magnitude is comparable to that from the continuum spectrum. The renormalization energies from the Feynman diagrams are dominated by $\Gamma_{2}$ which value is negative, while $\Gamma_{4}$ is positive and relatively small. By adding them all one obtains the total renormalized vacuum fluctuation energy over the chiral soliton background which is negative. If the color degrees of freedom of quarks are taken into account, the vacuum energy should be multiplied by a factor of 3. One could see that the quantum fluctuation energy in magnitude is not small compared to the classical energy of the chiral soliton.
@@ V SUMMARY AND OUTLOOK
Appendix A Renormalization by the fake boson method
In this appendix, we perform renormalization calculations about the logarithmic divergence using the fake boson method. We introduce the fake boson in a background potential
$V_{B}(r)=m_{B}^{2}\frac{r}{\rm w}e^{-3\frac{r}{\rm w}},$ (78)
where $\rm w$ is an arbitrary width which should not play a role in the final result, and $m_{B}$ is the mass of the fake boson. We consider the spherically symmetric problem
$\frac{d^{2}}{dr^{2}}u_{\ell}(r)+\left[\frac{\ell(\ell+1)}{r^{2}}+V_{B}(r)\right]u_{\ell}(r)=k^{2}u_{\ell}(r).$ (79)
In order to obtain the second order scattering phase shift with different angular quantum number $\ell$, we take the ansatz $u_{\ell}(r)=\exp(2i\beta_{\ell}(k,r))rh_{\ell}(kr)$ and the expansion $\beta_{\ell}(k,r)=g_{B}\beta_{\ell}^{(1)}(k,r)+g_{B}^{2}\beta_{\ell}^{(2)}(k,r)$, where $g_{B}$ is a formal coupling constant. Then the coupled differential equations about $\beta_{\ell}^{(1)}$ and $\beta_{\ell}^{(2)}$ could be derived as
$-i\beta_{\ell}^{(1)\prime\prime}-2ikp_{\ell}(kr)\beta_{\ell}^{(1)\prime}+\frac{1}{2}V_{B}(r)$ $=0,$ (80)
$-i\beta_{\ell}^{(2)\prime\prime}-2ikp_{\ell}(kr)\beta_{\ell}^{(2)\prime}+2\left(\beta_{\ell}^{(1)\prime}\right)^{2}$ $=0.$ (82)
By numerically solving the equations, one can obtain the Born subtraction phase shift of the second order as
$\delta_{\ell,B}^{(2)}(k)=2Re[\beta_{\ell}^{(2)}(k,0)].$ (83)
The logarithmic divergent energies from the second-order fake boson subtraction phase shift and from the third- and fourth-order fermion subtraction phase shifts could be evaluated in the following
$E_{B}^{(2)}(q)$ $=\int_{0}^{q}\frac{{\rm d}k}{\pi}\frac{d\delta_{B}^{(2)}(k)}{dk}\sqrt{k^{2}+m_{B}^{2}},$ (84)
$E_{F}^{(3,4)}(q)$ $=\int_{0}^{q}\frac{{\rm d}k}{\pi}\frac{d\delta_{F}^{(3,4)}(k)}{dk}\sqrt{k^{2}+m_{q}^{2}},$ (85)
where the Born phase shifts $\delta_{B}^{(2)}$ and $\delta_{F}^{(3,4)}$ are defined as
$\delta_{B}^{(2)}(k)$ $=2\sum_{\ell}(2\ell+1)\delta_{\ell,B}^{(2)}(k),$ (86)
$\delta_{F}^{(3,4)}(k)$ $=\sum_{G,\omega,\Pi}(2G+1)(\delta_{G,\omega,\Pi}^{(3)}(k)+\delta_{G,\omega,\Pi}^{(4)}(k)).$ (87)
Next we follow the strategy in Ref. Quandt et al. (2017) to match the renormalization energies of the fake boson and fermion diagrams. The one loop fake boson diagram related to the second order Born subtraction could be formally evaluated by a on-shell mass renormalization and the result is
$\Pi_{ren}(q)=-\int_{0}^{1}dx\frac{x(1-x)}{m_{B}^{2}-x(1-x)m_{f}^{2}}(q^{2}+m_{f}^{2})-\int_{0}^{1}dx\ln\frac{m_{B}^{2}+x(1-x)q^{2}}{m_{B}^{2}-x(1-x)m_{f}^{2}},$ (88)
where $m_{f}$ is a formal mass of the external line of the boson loop. The corresponding finite fake boson renormalization energy is
$\Gamma_{2B}=\int_{0}^{\infty}\frac{dqq^{2}}{2\pi^{2}}\tilde{V}_{B}(q)\tilde{V}_{B}(-q)\Pi_{ren}(q).$ (89)
The fake boson renormalization energy and the fermion renormalization energy could be matched by the following scaling constant
$c_{B}=\frac{\Gamma_{4}}{\Gamma_{2B}},\ \ \ c_{B}\equiv\lim_{q\to\infty}C_{B}(q)=\frac{E_{F}^{(3,4)}|_{q\to\infty}}{E_{B}^{(2)}|_{q\to\infty}},$ (90)
which can be numerically evaluated by the ratio of $E_{F}^{(3,4)}(q)$ and $E_{B}^{(2)}(q)$ when the momentum $q\to\infty$}. In the numerical evaluation the width parameter is set to ${\rm w}=2$, then the fake boson mass is determined at $m_{B}=2.5g\sigma_{v}$ to guarantee the scaling parameter approaching a constant when the momentum goes to infinity. The momentum dependent scaling parameter $C_{B}(q)$ is plotted in Fig.8. The numerical value of the scaling constant is fixed at $c_{B}=-14.564$. Thus one obtain the scaled renormalization energy which is equivalent to the renormalization energy of the fermion diagram
$\Gamma_{4}=c_{B}\,\Gamma_{2B}.$ (91)
However, in this renormalizaiton energy the mass parameter $m_{f}$ is unphysical and arbitrary. In the main text as we have adopted the renormalization scheme from Farhi’s work Farhi et al. (2002), the scaled renormalization energy calculated here should agree with the energy calculated according to the renormalization scheme in Farhi’s paper. In the chiral soliton model, by the Farhi’s scheme the renormalization energy could be written as
$\Gamma_{\rm loc}^{(4)}=$ $\frac{g^{4}}{8\pi}\left(\frac{m_{s}^{2}}{m_{q}^{2}}+6\int_{0}^{1}dx\ln\left[1-x(1-x)\frac{m_{s}^{2}}{m_{q}^{2}}\right]\right)$
$\times\int_{0}^{\infty}drr^{2}\left[\left(\sigma(r)^{2}+\pi(r)^{2}-\sigma_{v}^{2}\right)^{2}-4\sigma_{v}^{2}(\sigma(r)-\sigma_{v})^{2}\right],$ (92)
where $m_{s}$ and $m_{q}$ are the physical masses. The subscript denotes that the local approximation has been applied, which means the external momentum in the loop Feynman diagram is set to zero. Accordingly, our scaled renormalization energy at the local approximation could be written as
$\Gamma_{4,\rm loc}=\frac{c_{B}}{4\pi}\left(-\int_{0}^{1}dx\frac{x(1-x)m_{f}^{2}}{m_{B}^{2}-x(1-x)m_{f}^{2}}+\int_{0}^{1}dx\ln\left[1-x(1-x)\frac{m_{f}^{2}}{m_{B}^{2}}\right]\right)\int_{0}^{\infty}drr^{2}V_{B}(r)^{2}.$ (93)
The two energies, $\Gamma_{4,\rm loc}$ and $\Gamma_{\rm loc}^{(4)}$, must be equal. With the masses $m_{s}$ and $m_{q}$ set to their physical values, the mass parameter $m_{f}$ is subsequently determined to be $m_{f}=0.923g\sigma_{v}$. This yields a final value for the renormalization energy of $\Gamma_{4}=2.269{\rm fm}^{-1}$.
Figure 8: The ratio $C_{B}$ as a function of momentum $q$ at $\rm w=2$ and $m_{B}=2.5g\sigma_{v}$.
Appendix B Calculations of Born subtractions of phase shifts by the first order differential equations